# M3RS: Multi-robot, Multi-objective, and Multi-mode Routing and Scheduling

### v2

@@ III-C M3RS for Multi-Robot Disinfection
We consider a fleet of surface disinfection robots deployed in hospital-like environments to provide around-the-clock disinfection. Specifically, we use G-robots [32], which are far-UV-based mobile manipulators designed for surface disinfection. A disinfection mission is a collection of disinfection tasks spread in an environment such as a hospital. Each mission has a specified time horizon within which it should be completed. Further, the tasks must be done in their specified time window. The tasks can have different disinfection dose requirements. This dosage dictates disinfection quality, energy consumption, and the task service time. Each task can be disinfected in a particular set of modes, where the modes have different disinfection dosage options for that task.
@@
The disinfection dosage for UV-based surface disinfection is product of UV exposure duration and irradiance of the UV light source [1]. A higher disinfection dosage would require a longer exposure duration. Typically, the disinfection dosage is expressed in terms of percentage of population reduction. For example, a dosage of $D_{90}$ implies $90\%$ reduction of pathogenic population. A higher reduction percentage requires stronger disinfection dose which in turn leads to longer exposure times. Longer exposure times further implies higher energy consumption. We consider four disinfection modes that are $D_{99.9999}$, $D_{99.99}$, $D_{99}$, and $D_{90}$. Here, $D_{99.9999}$ is the strongest disinfection dose with the highest disinfection time and energy consumption as compared to $D_{90}$ [1]. The set of modes can be determined by the acceptable disinfection levels for a surface. So, each task has an associated set of acceptable disinfect doses or modes. The assignment of robots to different tasks and the corresponding modes will be determined by solving M3RS.
@@
For disinfection applications, we have two objectives: i) the number of tasks completed and ii) overall disinfection quality. Higher disinfection quality requires more time and resources, limiting availability for other tasks. For example, critical locations like ICUs in hospitals require higher dosages, so higher quality is preferable. In contrast, areas like waiting rooms or offices do not need the highest quality, and covering the maximum number of surfaces might be preferable. Our formulation creates schedules by using user-specified trade-off requirements. The formulation of M3RS for disinfection application is:
$\underset{x^{k}_{i{\xrightarrow{}}j},y^{k}_{i,m}}{\max}\,\frac{1}{|\mathcal{T}|}\cdot\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}(\lambda)\cdot x^{k}_{i{\xrightarrow{}}j}+(1-\lambda)\cdot\frac{p_{i,m}}{p^{max}}\cdot y^{k}_{i,m,}$ (23)
subject to constrains in Eq. (3)-(16). Here, $p_{i,m}$ indicates the quality of disinfection when completing task $i$ in mode $m$, $p^{max}$ is the highest possible quality of disinfection, and $\lambda$ represents the user-specified preference. When $\lambda=1$, the objective ignores quality and focuses on maximizing the number of tasks completed, whereas when $\lambda=0$, the objective exclusively prioritizes maximizing the quality of disinfection. This mixed integer optimization model can be solved using a standard branch and cut algorithm. However, for larger problem instances, the branch and cut solution has an exploding computation time. The CCG algorithm described in Alg. 1 can be used to effectively compute solutions in such cases.
(a) A sample layout for synthetic experiment.
(b) Comparison of idling time to $T_{H}$ ratio for M3RS and RS-F for different number of tasks and a fleet size of 4.
(c) Comparison of working time to $T_{H}$ ratio for M3RS and RS-F for different number of tasks and a fleet size of 4.
(d) Impact on $SR$ and $DQ$ for different types of time windows. The tasks with time-windows could be critical or non-critical.
Figure 2: (a) A sample layout of a disinfection mission with different task types. (b) & (c) The solution of M3RS make effective use of agents as indicated by lower idling times and higher working times. (d) It can be seen that tasks available during the entire mission have a higher completion rate ($SR$) and disinfection quality as well.
@@ IV Evaluation
In this section, we evaluate the the M3RS problem and propesed solution strategy of CCG. First, the metrics used for evaluation are defined. Second, the impact of using task-level quality as a decision variable is studied by comparing M3RS with the classic routing and scheduling in fixed (RS-F) disinfection mode on synthetically generated data. Then, the proposed CCG scheme is compared with a mixed integer programming (MIP) solver. Finally, hardware and simulated case studies are presented.
In each evaluation, multiple problem instances are generated. Each problem instance is characterized by mission time which is given by:
$T_{H}=\frac{t_{avg}\times\text{\# Tasks}}{\eta\times\text{Fleet size}}$ (24)
Here, $\eta$ is the load factor that is used to simulate missions with different $T_{H}$. A higher $\eta$ leads to tighter mission times. The average task completion time is given by $t_{avg}$. In our experiments, $t_{avg}$ is considered to be 4.5 minutes. We assume that the robot travels at a fixed speed of $0.5\,{m}/{s}$. We use the MIP solver from IBM’s CPLEX [33]. All the evaluations are carried out on machine equipped with an Intel i9 processor with 32 GB RAM.
@@ IV-A Evaluation Metrics
We define some evaluation metrics to compare the performance of M3RS and RS-F (Sec. IV-B). The metrics are:
$SR\,=\,\frac{\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}j}}{|\mathcal{T}|}$ (25)
$DQ\,=\,\frac{\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}p_{i,m}\cdot y^{k}_{i,m}}{|\mathcal{T}|\cdot p^{max}}$ (26)
$MSI\,=\,\frac{SR+DQ}{2}$ (27)
$ATQ\,=\,\frac{DQ}{SR}$ (28)
$QBR\,=\,\frac{DQ-DQ^{*}}{(\bar{Q}_{exp}-\bar{Q}_{exp}^{*})+(\bar{M}_{exp}-\bar{M}_{exp}^{*})}$ (29)
@@
The quality benefit ratio ($QBR$) compares a target solution with a baseline solution in terms of the gains made in disinfection quality relative to the resource usage (Eq. (29)). It is a ratio of the difference in disinfection quality to the difference in resource expenditure, measured against a baseline solution. For the target solution, $\bar{Q}_{exp}$ and $\bar{M}_{exp}$ represent the fleet’s average energy expenditure and makespan. The RS-F solution serves as the baseline with $DQ^{*}$ as disinfection quality, $\bar{Q}_{exp}^{*}$ as average energy, and $\bar{M}_{exp}^{*}$ as makespan.
In Sec IV-C, we use relative optimal gap ($OG$) and the objective values (Eq. (23)) are also as a metric. The $OG$ measures the difference between the best known bound on the objective value ($o^{*}$) and the current solution’s objective value $o$, is also reported. MIP solvers typically provide this bound by solving a relaxed version of the problem and is given by:
$OG\,=\,\frac{o^{*}-o}{o^{*}}$ (30)
IV-B M3RS vs RS-F
IV-B1 Baselines
We assess the impact of the M3RS formulation, which incorporates quality levels as a decision variable through modes, compared to RS-F which is routing and scheduling with fixed execution modes. The RS-F formulation follows the vehicle routing problem with time windows [34], and we consider two variants: i) RS-F (Min), which uses the mode with the lowest disinfection quality, and ii) RS-F (Max), which uses the mode with the highest quality. For M3RS we use different values of user prefrence i.e. $\lambda$. A higher $\lambda$ prioritizes quantity over quality. Synthetically generated data is used to benchmark the problems, as detailed in the following paragraph. Both M3RS and RS-F (Min and Max) are solved using an MIP solver with a 10-minute time limit per problem instance. The evaluation metrics include $SR$, $DQ$, $MSI$ $ATQ$, and $QBR$, see Eq (25)-(29). RS-F (Min) is used as the baseline for $QBR$ calculation since it operates in the lowest disinfection mode, allowing us to measure the improvement in disinfection quality relative to resource expenditure. We analyze the solutions based on working and idle time and also discuss the impact of relaxing task time windows for M3RS.
IV-B2 Synthetic Data Generation
[…]
Table II: Experiment parameters for testing impact of quality.
@@ IV-B2 Synthetic Data Generation
Avg. task completion time $4.5\,mins$
@@
Table III: Comparison of performance metrics for synthetic problem instances. The colors indicate the first, second, and third ranking for a given metric. The solutions of M3RS perform well on both quality and quantity objectives as indicated by different metrics. The number of tasks range from $30-60$ and the robot fleet size is four.
Metric RS-F (Min) RS-F (Max) M3RS-0.01 M3RS-0.1 M3RS-0.5 M3RS-0.9
$SR$ $1.0\pm 0.0$ $0.847\pm 0.082$ $0.906\pm 0.078$ $0.912\pm 0.072$ $0.941\pm 0.068$ $0.973\pm 0.042$
$DQ$ $0.452\pm 0.031$ $0.757\pm 0.083$ $0.820\pm 0.078$ $0.829\pm 0.097$ $0.789\pm 0.095$ $0.750\pm 0.124$
$MSI$ $0.726\pm 0.016$ $0.802\pm 0.082$ $0.863\pm 0.072$ $0.870\pm 0.079$ $0.865\pm 0.073$ $0.862\pm 0.08$
$ATQ$ $0.452\pm 0.031$ $0.894\pm 0.024$ $0.906\pm 0.063$ $0.909\pm 0.070$ $0.839\pm 0.082$ $0.768\pm 0.104$
$QBR$ N/A $0.331\pm 0.064$ $0.364\pm 0.082$ $0.391\pm 0.117$ $0.331\pm 0.104$ $0.297\pm 0.129$
Table IV: Comparison of $OG$ values between MIP solver and CG solutions. Colors represent the first, second, and third rankings for the $OG$ metric. CCG solutions achieve better $OG$ values as compared to the MIP solver for the same computation time limit.
Method Number of Tasks and Robots, $\eta=1.0$, $\lambda=0.9$ Number of Tasks and Robots, $\eta=2.0$, $\lambda=0.9$
120-4 150-5 175-5 200-6 120-4 150-5 175-5 200-6
MIP $0.131\pm 0.034$ $0.278\pm 0.113$ $0.247\pm 0.046$ $0.353\pm 0.024$ $0.253\pm 0.032$ $0.290\pm 0.061$ $0.570\pm 0.065$ $0.488\pm 0.039$
CCG ($K=2$) $0.136\pm 0.050$ $0.110\pm 0.024$ $0.218\pm 0.024$ $0.300\pm 0.018$ $0.116\pm 0.015$ $0.177\pm 0.039$ $0.455\pm 0.011$ $0.380\pm 0.006$
CCG ($K=4$) $0.126\pm 0.009$ $0.121\pm 0.034$ $0.211\pm 0.017$ $0.302\pm 0.023$ $0.097\pm 0.045$ $0.179\pm 0.021$ $0.434\pm 0.013$ $0.363\pm 0.014$
CCG ($K=6$) $0.110\pm 0.025$ $0.132\pm 0.025$ $0.208\pm 0.010$ $0.305\pm 0.021$ $0.107\pm 0.025$ $0.180\pm 0.021$ $0.453\pm 0.026$ $0.358\pm 0.029$
CCG ($K=8$) $0.133\pm 0.031$ $0.102\pm 0.022$ $0.220\pm 0.019$ $0.308\pm 0.009$ $0.106\pm 0.047$ $0.184\pm 0.020$ $0.446\pm 0.019$ $0.355\pm 0.025$
Method Number of Tasks and Robots, $\eta=1.0$, $\lambda=0.1$ Number of Tasks and Robots, $\eta=2.0$, $\lambda=0.1$
120-4 150-5 175-5 200-6 120-4 150-5 175-5 200-6
MIP $0.124\pm 0.032$ $0.181\pm 0.021$ $0.153\pm 0.014$ $0.215\pm 0.019$ $0.409\pm 0.031$ $0.586\pm 0.013$ $0.5200\pm 0.026$ $0.602\pm 0.021$
CCG ($K=2$) $0.103\pm 0.018$ $0.144\pm 0.037$ $0.188\pm 0.022$ $0.205\pm 0.016$ $0.357\pm 0.037$ $0.483\pm 0.015$ $0.425\pm 0.018$ $0.514\pm 0.021$
CCG ($K=4$) $0.116\pm 0.018$ $0.126\pm 0.019$ $0.156\pm 0.029$ $0.197\pm 0.020$ $0.310\pm 0.023$ $0.491\pm 0.016$ $0.436\pm 0.021$ $0.517\pm 0.018$
CCG ($K=6$) $0.132\pm 0.027$ $0.150\pm 0.034$ $0.162\pm 0.032$ $0.201\pm 0.020$ $0.349\pm 0.031$ $0.483\pm 0.012$ $0.418\pm 0.009$ $0.520\pm 0.014$
CCG ($K=8$) $0.116\pm 0.019$ $0.134\pm 0.015$ $0.198\pm 0.010$ $0.181\pm 0.027$ $0.336\pm 0.011$ $0.491\pm 0.012$ $0.424\pm 0.015$ $0.502\pm 0.025$

### v3

@@ III-C M3RS for Multi-Robot Disinfection
We consider a fleet of surface disinfection robots deployed in hospital-like environments to provide around-the-clock disinfection. Specifically, we use G-robots (Mehta et al.,, 2022), which are mobile manipulators equipped with far-UVC emitters and are designed for surface disinfection (see Fig. 7a). A disinfection mission is a collection of disinfection tasks spread in an environment such as a hospital. Each mission has a specified time horizon within which it should be completed. Further, the tasks must be done in their specified time window. The tasks can have different disinfection dose requirements. This dosage dictates disinfection quality, energy consumption, and the task service time. Each task can be disinfected in a particular set of modes, where the modes have different disinfection dosage options for that task.
@@
The disinfection dosage ($D$ in $\frac{\text{Watt}\cdot s}{m^{2}}$) for UV-based surface disinfection is the product of UV exposure duration ($E_{t}$ in $s$) and irradiance of the UV light source ($I_{r}$ in $\frac{\text{Watt}}{m^{2}}$), given by $D=E_{t}\cdot I_{r}$ (Mehta et al.,, 2023). A higher disinfection dosage ($D$), requires a longer exposure time ($E_{t}$), given the fixed irradiance of the UV light source ($I_{r}$). Disinfection dosage is typically expressed as the percentage reduction in pathogen populations and is pre-calculated for various pathogen types (Mehta et al.,, 2023). For example, a dosage of $D_{90}$ implies a $90\%$ reduction of the pathogenic population. A higher reduction percentage requires a stronger disinfection dose, which in turn leads to longer exposure times. Longer exposure times further imply higher energy consumption. We consider four disinfection modes that are $D_{99.9999}$, $D_{99.99}$, $D_{99}$, and $D_{90}$. Here, $D_{99.9999}$ is the strongest disinfection dose with the highest disinfection time and energy consumption as compared to $D_{90}$ (Mehta et al.,, 2023). The set of modes can be determined by the acceptable disinfection levels for a surface. So, each task has an associated set of acceptable disinfectant doses or modes. The assignment of robots to different tasks and the corresponding modes will be determined by solving M3RS. The multi-objective formulation and execution mode decision variables in M3RS enable trade-offs between disinfection quality and number of tasks completed, based on mission constraints and user preferences.
@@
For disinfection applications, we have two objectives: i) the number of tasks completed and, i.e. $f_{1}$, ii) overall disinfection quality i.e. $f_{2}$. Higher disinfection quality requires more time and resources, limiting availability for other tasks. For example, critical locations like ICUs in hospitals require higher dosages, so higher quality is preferable. In contrast, areas like waiting rooms or offices do not need the highest quality, and covering the maximum number of surfaces might be preferable. Our formulation creates schedules by using user-specified trade-off requirements. The overall objective of M3RS–composed of $f_{r}(\cdot),\penalty 10000\ r=1,2$, in Eq. (1)–for the disinfection application is given by:
$\underset{x^{k}_{i{\xrightarrow{}}j},y^{k}_{i,m}}{\max}\,\frac{1}{|\mathcal{T}|}\cdot\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}(\lambda)\cdot f_{1}+(1-\lambda)\cdot f_{2},$ (23)
subject to the constraints in Eqs. (2)–(16). This $f_{1}(\cdot)$ and $f_{2}(\cdot)$ are defined as follows:
$f_{1}=x^{k}_{i\rightarrow j},$ (24)
which maximizes the quantity of tasks completed, and
$f_{2}=\frac{p_{i,m}}{p^{\max}}\cdot y^{k}_{i,m},$ (25)
which maximizes the quality of completed tasks. Here, $p_{i,m}$ denotes the reward representing the disinfection quality for task $i$ executed in mode $m$ (defined in Sec. IV-B2), $p^{\text{max}}$ is the maximum achievable disinfection quality reward, and $\lambda$ captures the user-defined preference. See Table I for a summary of the notation. When $\lambda=1$, the objective prioritizes maximizing the number of tasks completed, disregarding disinfection quality. Conversely, when $\lambda=0$, it exclusively focuses on maximizing disinfection quality. Users can adjust $\lambda$ based on mission priorities. Lower values suit missions with critical areas like ICUs or patient care rooms. Higher values are appropriate for missions with less critical spaces, e.g., waiting rooms. This mixed integer optimization model can be solved using a standard branch-and-cut algorithm. However, for larger problem instances, the branch-and-cut has an exponentially increasing computation time. The CCG algorithm, Alg. 1, can be used in such cases.
@@ IV Evaluation
In this section, the M3RS problem and the proposed solution of CCG are evaluated. The evaluation is conducted for multi-robot disinfection scenarios. First, the metrics used for evaluation are defined. Then, the impact of using task-level disinfection quality as a decision variable is analyzed. This is done by comparing M3RS with a baseline that performs classic routing and scheduling in fixed (RS-F) disinfection modes. Fixed modes are used since existing MRTA methods do not treat quality as a decision variable, and no relevant baselines exist. The comparison is performed on synthetically generated multi-robot disinfection missions. Additioannly, the proposed CCG scheme is compared with a mixed integer programming (MIP) solver on synthetically generated missions. Finally, hardware and simulated case studies are presented.
In each evaluation, multiple problem instances are generated. Each instance is characterized by mission time, given by:
$T_{H}=\frac{t_{avg}\times\text{\# Tasks}}{\eta\times\text{Fleet size}}.$ (26)
Here, $T_{H}$ represents the mission time, i.e., the total time available to complete all tasks, while $\eta$ is the load factor used to simulate missions with varying $T_{H}$. The load factor is commonly used in MRTA to represent mission workload relative to the robot’s available capacity (Botros et al.,, 2023). We use the load factor to model varying mission durations, where a higher $\eta$ corresponds to tighter time constraints for the mission. The average task completion time is given by $t_{avg}$. In our experiments, $t_{avg}$ is considered to be 4.5 minutes. We assume that each robot travels at a fixed speed of 0.5m/s0.5\,{m}/{s}. We use the MIP solver from IBM’s CPLEX (Laborie et al.,, 2018). All evaluations are done on a machine with an Intel i9 processor and 32GB RAM.
@@ IV-A Evaluation Metrics
We define six evaluation metrics to compare the performance of M3RS against other methods. The metrics are:
$SR\,=\,\frac{\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}j}}{|\mathcal{T}|},$ (27)
$DQ\,=\,\frac{\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}p_{i,m}\cdot y^{k}_{i,m}}{|\mathcal{T}|\cdot p^{max}},$ (28)
$MSI\,=\,\frac{SR+DQ}{2},$ (29)
$ATQ\,=\,\frac{DQ}{SR},$ (30)
$QBR\,=\,\frac{DQ-DQ^{*}}{(\bar{Q}_{exp}-\bar{Q}_{exp}^{*})+(\bar{M}_{exp}-\bar{M}_{exp}^{*})}.$ (31)
@@
The quality-benefit ratio ($QBR$) compares a target solution with a baseline solution in terms of the gains made in disinfection quality relative to the resource usage (Eq. (31)). It is a ratio of the difference in disinfection quality to the difference in resource expenditure, measured against a baseline solution. For the target solution, $\bar{Q}_{exp}$ and $\bar{M}_{exp}$ represent the fleet’s average energy expenditure and makespan. The RS-F solution serves as the baseline with $DQ^{*}$ as disinfection quality, $\bar{Q}_{exp}^{*}$ as average energy, and $\bar{M}_{exp}^{*}$ as makespan, where makespan is time taken by robot to complete its assigned task during the mission.
The objective values (Eq. (23)) are also used as a metric. This is achieved by measuring the difference between the best-known bound on the objective value ($o^{*}$) and the current solution’s objective value, $o$. MIP solvers typically provide the bound by solving a relaxed version of the problem. This metric is presented as the relative optimality gap ($OG$) (Kaliszewski and Miroforidis,, 2021):
$OG\,=\,\frac{o^{*}-o}{o^{*}}\cdot 100\,\%.$ (32)
@@
Avg. task completion time 4.5mins4.5\,mins
@@
$Q$ $6\text{Ah}$
Table II: A list of parameters used to simulate disinfection missions to assess the impact of quality. A mission consists of disinfection tasks spread across a space, see Fig. 3.
Figure 3: A sample physical layout of locations of different disinfection tasks. Here, the color coding is used to present different categories of tasks.
IV-B M3RS vs RS-F (Routing and Scheduling with Fixed modes)
In this experiment, we assess the effectiveness of our proposed M3RS formulation in multi-robot disinfection scenarios. The results demonstrate that M3RS outperforms baseline approaches across key performance metrics.
IV-B1 Baselines
We evaluate the impact of the M3RS formulation, which integrates quality levels as a decision variable through multiple execution modes, against RS-F, a routing and scheduling approach with fixed execution modes. RS-F follows the vehicle routing problem with time windows (Kallehauge et al.,, 2005) and does not account for task-level quality. We consider two RS-F variants: i) RS-F (Min), which defaults to the lowest disinfection dosage, and ii) RS-F (Max), which defaults to the highest disinfection dosage. For M3RS we use different values of user preference i.e. $\lambda$. A higher $\lambda$ prioritizes quantity over quality. Synthetically generated data is used to benchmark the problems, as detailed in the following paragraph. Both M3RS and RS-F (Min and Max) are solved using an MIP solver with a 10-minute time limit per problem instance. The evaluation metrics include $SR$, DQDQ, MSIMSI ATQATQ, and $QBR$, see Eq. (27)-(31). RS-F (Min) is used as the baseline for $QBR$ calculation since it operates in the lowest disinfection mode, allowing us to measure the improvement in disinfection quality relative to resource expenditure. We analyze the solutions based on working and idle time and also discuss the impact of relaxing task time windows for M3RS.
IV-B2 Synthetic Data Generation
The synthetic data simulates realistic hospital missions with rectangular layouts of varying sizes. For each layout type, $10$ random problem instances are generated and solved. A sample layout is shown in Fig. 3. We consider four disinfection modes, indexed by $m$, where a lower $m$ indicates higher disinfection quality. Specifically, $m=0$ corresponds to $D_{99.9999}$ (highest quality), followed by $D_{99.99}$ ($m=1$), $D_{99}$ ($m=2$), and $D_{90}$ ($m=3$, lowest quality). For task $i$, the mode set ($\mathcal{M}_{i}$) is randomly assigned between one and four disinfection modes. The associated quality reward (see Eq. (23)) for each mode is defined as $p_{i,m}=2^{-m}$ for $m\in\{0,1,2,3\}$. For example, selecting the highest quality mode ($m=0$) for task $i$ yields the maximum reward of $1$, while the lowest quality mode ($m=3$) yields a reward of $0.125$. For each disinfection task, the target surface is assumed to be planar, with an area randomly assigned between $0.5$ and $2.2\,m^{2}$. The required dosage, exposure time, and energy consumption for disinfecting the surface are computed based on UV disinfection dosage and energy models (Mehta et al.,, 2022). A summary of the experimental setup is provided in Table II. Each problem instance involves a fleet of four robots. We set the value of load factor $\eta=1$ for this experiment. This ensures that for the disinfection mission there is sufficient time to address the tasks. The mission time ($T_{H}$) is calculated using Eq. (26). Tasks are categorized as critical or non-critical. In each problem instance, a subset of critical tasks, representing $15-25\%$ of the total tasks, must be completed within a specified time window using the highest disinfection dosage, $D_{99.9999}$. These are crucial surfaces, such as those in patient rooms and ICUs. The inequality in Eq. (2) is replaced by an equality constraint for critical tasks. Non-critical tasks, like those in waiting areas, washrooms, and office spaces, can be skipped but contribute to overall hygiene if addressed. These tasks may have fixed time windows, and deadlines, or be available throughout the mission. All time windows are uniformly sampled from the mission time horizon.